What velocity should a clock be moved related to the observer so that it may seems to lose 1 minute per day
1
Expert's answer
2019-05-09T11:20:56-0400

Time in relativistic physics is determined as

"T=\frac{T_0}{\sqrt{1-\frac{v^2}{c^2}}}"

So, if The interval for observer

"T_0=86400 (s)"

then interval for clock be moved related to the observer

"T=86460 (s)"

Velocity is

"v=c\sqrt{1-(\frac{T_0}{T})^2}"

calculating

"v=3\cdot 10^8\cdot \sqrt{1-(\frac{86400}{86460})^2}=11.2 (Mm\cdot s^{-1})"